# Linear complementary dual quasi-cyclic codes of index 2

### v1

@@ 6 Hermitian LCD quasi-cyclic codes of index 2
Recall that the conjugate of a polynomial $f(x)=f_{0}+f_{1}x+\dots+f_{k}x^{k}\in\mathbb{F}_{q^{2}}[x]$ of degree $k$ is defined as
@@
The map ^$\hat{\ }:f(x)\mapsto\hat{f}(x)$ can be naturally extended to the following isomorphisms:
^$\begin{split}\hat{\ }:\ \mathbb{F}_{q^{2}}[x]/(f_{i}(x))&\to\mathbb{F}_{q^{2}}[x]/(f_{i}(x)),\\ \hat{\ }:\ \mathbb{F}_{q^{2}}[x]/(h_{j}(x))&\to\mathbb{F}_{q^{2}}[x]/(h_{j}^{\dagger}(x)).\\ \end{split}$ (8)
(Here ^$\hat{\ }:a(x)+(h_{j}(x))\mapsto\hat{a}(x)+(h_{j}^{\dagger}(x))$.) Therefore, the map $\hat{\ }$ is an isomorphism from $H_{j}^{\prime}=\mathbb{F}_{q^{2}}[x]/(h_{j}(x))$ to $H_{j}^{\prime\prime}=\mathbb{F}_{q^{2}}[x]/(h_{j}^{\dagger}(x))$.
@@
${H^{\prime}_{j}}$${\mathbb{F}_{q^{2}}(\xi^{v_{j}})}$${H^{\prime\prime}_{j}}$$\scriptstyle{\mu_{j}}$$\scriptstyle{\widehat{\ }}$$\scriptstyle{\nu_{j}}$
@@
The constituent code $C_{j}^{\prime}$ over $H_{j}^{\prime}$ is obtained by evaluating $a(x)$ at $\xi^{v_{j}}$, while the constituent code $C_{j}^{\prime\prime}$ over $H_{j}^{\prime\prime}$ is obtained by evaluating $\hat{a}(x)$ at $\xi^{v_{j}}$. In [4], the authors utilized a slightly different isomorphism. They defined the constituent code $C_{j}^{\prime}$ over $H_{j}^{\prime}$ by evaluating $a(x)$ at $\xi^{v_{j}}$ and the constituent code $C_{j}^{\prime\prime}$ over $H_{j}^{\prime\prime}$ by evaluating $a(x)$ at $\xi^{-qv_{j}}$, which means $a(\xi^{-qv_{j}})=(\hat{a}(\xi^{v_{j}}))^{q}$ (see [4, Eq. (IV.10)] and remark after Eq. (IV.14)).
@@
We have the following characterization of Hermitian LCD QC codes in terms of constituents (see [4, Eq. (IV.16)]).
@@
Let $C$ be a QC code of length $2m$ and index $2$ over $\mathbb{F}_{q^{2}}$, generated by $(g_{11}(x),g_{12}(x))$ and $(0,g_{22}(x))$ satisfying Conditions ()($\ast$)( ∗ ). Similar to Section 3, in the above setting, each constituent of $C$ is generated by the rows of a $2\times 2$ matrix over its field of definition. Explicitly, $C_{i},C_{j}^{\prime}$ and $C_{j}^{\prime\prime}$ are generated by the rows of the matrices
@@
Let $C$ be a quasi-cyclic code of index $2$. Let $(g_{11}(x),g_{12}(x))$ and $(0,g_{22}(x))$ be the generators of $C$ satisfying Conditions ()($\ast$)( ∗ ). Then $C$ is Hermitian LCD if and only if all of the following conditions are true:
@@ Conclusion
In this work, we have given a nice polynomial-based characterization of quasi-cyclic linear complementary dual (LCD) codes of index $2$ with respect to the Euclidean, Hermitian and symplectic inner products. Our results extend the existing characterizations obtained for one-generator quasi-cyclic codes. Moreover, the techniques introduced in our characterization can be readily generalized to the broader class of quasi-twisted codes. A promising direction for future research is to obtain polynomial characterization for quasi-cyclic LCD codes of arbitrary index. This extension would require handling a large number of polynomials together, which will be more complicated.
@@ Acknowledgment
K. Abdukhalikov was supported by UAEU grants G00004233 and G00004614. D. Ho was supported by the Tromsø Research Foundation (project “Pure Mathematics in Norway”) and UiT Aurora project MASCOT. G. K. Verma was supported by UAEU grant G00004614. S. Ling is supported by Nanyang Technological University Research Grant No. 04INS000047C230GRT01 and UAEU grant G00004233.

### v2

@@ 6 Hermitian LCD quasi-cyclic codes of index 2
Recall that the conjugate of a polynomial $f(x)=f_{0}+f_{1}x+\dots+f_{k}x^{k}\in\mathbb{F}_{q^{2}}[x]$ is defined as
@@
The map $\hat{\ }:f(x)\mapsto\hat{f}(x)$ can be naturally extended to the following isomorphisms:
[…]
(Here $\hat{\ }:a(x)+(h_{j}(x))\mapsto\hat{a}(x)+(h_{j}^{\dagger}(x))$.) Therefore, the map $\hat{\ }$ is an isomorphism from $H_{j}^{\prime}=\mathbb{F}_{q^{2}}[x]/(h_{j}(x))$ to $H_{j}^{\prime\prime}=\mathbb{F}_{q^{2}}[x]/(h_{j}^{\dagger}(x))$.
@@
$H^{\prime}_{j}$$\mathbb{F}_{q^{2}}(\xi^{v_{j}})$$H^{\prime\prime}_{j}$$\scriptstyle{\mu_{j}}^\scriptstyle{\widehat{\ }}$$\scriptstyle{\nu_{j}}$
@@
where each component codes $C_{i}$, $C_{j}^{\prime}$ and $C_{j}^{\prime\prime}$ are constituent linear codes of length $\ell$ over the fields $F_{i}$, $H_{j}^{\prime}$, and $H_{j}^{\prime\prime}$, respectively.
@@
The constituent code $C_{j}^{\prime}$ over $H_{j}^{\prime}$ is obtained by evaluating $a(x)$ at $\xi^{v_{j}}$, while the constituent code $C_{j}^{\prime\prime}$ over $H_{j}^{\prime\prime}$ is obtained by evaluating $\hat{a}(x)$ at $\xi^{v_{j}}$. In [11], the authors utilized a slightly different isomorphism. They defined the constituent code $C_{j}^{\prime}$ over $H_{j}^{\prime}$ by evaluating $a(x)$ at $\xi^{v_{j}}$ and the constituent code $C_{j}^{\prime\prime}$ over $H_{j}^{\prime\prime}$ by evaluating $a(x)$ at $\xi^{-qv_{j}}$, which means $a(\xi^{-qv_{j}})=(\hat{a}(\xi^{v_{j}}))^{q}$ (see [11, Eq. (IV.10)] and remark after Eq. (IV.14)).
@@
We have the following characterization of Hermitian LCD QC codes in terms of constituents (see [11, Eq. (IV.16)]).
@@
Let $C$ be a QC code of length $2m$ and index $2$ over $\mathbb{F}_{q^{2}}$, generated by $(g_{11}(x),g_{12}(x))$ and $(0,g_{22}(x))$ satisfying Conditions $(\ast)$. Similar to Section 3, in the above setting, each constituent of $C$ is generated by the rows of a $2\times 2$ matrix over its field of definition. Explicitly, $C_{i},C_{j}^{\prime}$ and $C_{j}^{\prime\prime}$ are generated by the rows of the matrices
@@
Let $C$ be a quasi-cyclic code of index $2$. Let $(g_{11}(x),g_{12}(x))$ and $(0,g_{22}(x))$ be the generators of $C$ satisfying Conditions $(\ast)$. Then $C$ is Hermitian LCD if and only if all of the following conditions are true:
@@
Next, using Theorem 6.3, we construct several new and optimal quasi-cyclic Hermitian LCD codes over $\mathbb{F}_{4}=\mathbb{F}_{2}[w]$ with $w^{2}+w+1=0$ (in Table 2).
$C$ $g_{11}(x)$ $g_{12}(x)$ $g_{22}(x)$ Remark
$[22,11,8]$ $x+1$ $(x+1)$ $(w^{2}x+wx^{2}$ $x^{4}+wx^{6})$ $(x^{5}+wx^{4}+x^{3}$ $x^{2}+w^{2}x+1)$ $(x^{5}+w^{2}x^{4}+x^{3}$ $x^{2}+wx+1)$ BKLC
$[22,16,4]$ $x+1$ $x^{5}+wx^{4}$ $x^{2}+x+1$ $x^{5}+wx^{4}+x^{3}$ $x^{2}+w^{2}x+1$ Optimal Hermitian LCD
$[26,13,8]$ $x+1$ $w^{2}x^{9}+x^{8}+w^{2}x^{7}$ $x^{6}+x^{4}+wx^{3} +$$x^{2}+wx$ $(x^{6}+wx^{5}+w^{2}x^{3}+wx+1)$ $(x^{6}+w^{2}x^{5}+wx^{3}+w^{2}x+1)$ -
$[34,16,11]$ $(x+1)(x^{4}+wx^{3}$ $x^{2}+wx+1)$ $(x+1)$ $(x^{9}+w^{2}x^{8}+x^{7}$ $wx^{2}+w^{2}x)$ $(x+1)$ $(x^{4}+x^{3}+wx^{2}+x+1)$ $(x^{4}+x^{3}+w^{2}x^{2}+x+1)$ $(x^{4}+w^{2}x^{3}+x^{2}+w^{2}x+1)$ New
$[34,17,9]$ $(x+1)(x^{4}+w^{2}x^{3}$ $x^{2}+w^{2}x+1)$ $x^{10}+x^{8}+w^{2}x^{7}$ $w^{2}x^{6}+x^{2}+x$ $(x^{4}+x^{3}+wx^{2}+x+1)$ $(x^{4}+x^{3}+w^{2}x^{2}+x+1)$ $(x^{4}+wx^{3}+x^{2}+wx+1)$ -
Table 2: Quaternary quasi-cyclic (Hermitian) LCD codes, where BKLC means that the minimum distance of the code is equal to the minimum distance of Best Known Linear Code according to the Grassl’s table, “New” means codes are new compare to papers [14],[20],[42] and “Optimal Hermitian LCD” means there exists no LCD code with parameters $[22,16,5]$ (see [31]).
@@ Conclusion
In this work, we have given a nice polynomial-based characterization of quasi-cyclic linear complementary dual (LCD) codes of index $2$ with respect to the Euclidean, Hermitian and symplectic inner products. Using these characterizations, we constructed several binary, ternary, and quaternary LCD codes. Our results extend the existing characterizations obtained for one-generator quasi-cyclic codes. Moreover, the techniques introduced in our characterization can be readily generalized to the broader class of quasi-twisted codes. A promising direction for future research is to obtain polynomial characterization for quasi-cyclic LCD codes of arbitrary index. This extension would require handling a large number of polynomials together, which will be more complicated.
@@ Acknowledgment
K. Abdukhalikov was supported by UAEU grants G00004233 and G00004614. D. Ho was supported by the Tromsø Research Foundation (project “Pure Mathematics in Norway”) and UiT Aurora project MASCOT. G. K. Verma was supported by UAEU grant G00004614. S. Ling was supported by Nanyang Technological University Research Grant No. 04INS000047C230GRT01 and UAEU grant G00004233.